### v1

# 1]Univ. Artois, UR 4025, Laboratoire Systèmes Electrotechniques et Environnement (LSEE), F-62400 Béthune, France 2]Univ. Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels based transforms: pooling operators and input representations matter

Abstract
This article presents a new approach based on MiniRocket, called SelF-Rocket, for fast time series classification (TSC). Unlike existing approaches based on random convolution kernels, it dynamically selects the best couple of input representations and pooling operator during the training process. SelF-Rocket achieves state-of-the-art accuracy on the University of California Riverside (UCR) TSC benchmark datasets.
keywords:
Time series classification, ROCKET, MiniRocket, Pooling operators, Input representation, Feature selection.1 Introduction
Random convolution kernels based transforms, ROCKET [1] and its successors, MiniRocket [2], MultiRocket [3] and others, have revolutionized the field of time series classification, offering an unrivaled compromise between speed and accuracy [4].
The main idea behind ROCKET methods is to use a very large number of random convolution kernels to transform the time series and apply one or more pooling operators to generate a set of features. A simple linear classifier can then be used.
To the best of our knowledge, all the proposed transformations rely on one or more fixed pooling operators such as, for example, PPV (Proportion of positive values), for feature extraction from time series.
In this article, a new approach based on MiniRocket, called SelF-Rocket, is developed. Unlike existing ones based on random convolution kernels, it dynamically selects the best data input representations and pooling operator combination during the training process.
This article is organized as follows. The fundamentals of time series classification and the main random convolution kernels based methods are recalled in Section 2. A statistical analysis, conducted on the University of California Riverside (UCR) archive [5], is then presented in Section 3 to motivate this work. It shows that while PPV is a pooling operator of choice, in most cases it is not the best one. The method SelF-Rocket is then introduced in Section 4 and its performances on the UCR archive are analyzed in Section 5. Finally, Section 6 offers a conclusion with some perspectives.
2 Related works
2.1 Time series classification
Time series classification (TSC) is a type of supervised learning problem where the objective is to assign a class label to a given time series instance. A Time series data consist of sequences of observations collected at constant intervals over time. Formally, a time series can be represented as:
where denotes the value of the time series at time step , and represents the length of the time series. This notation will be used to refer to a time series throughout this article.
Many methods of various kind have been developed to address TSC problems. Traditional methods include distance-based techniques like Dynamic Time Warping (DTW) [6], which measure the similarity between time series. More recently, Deep Learning approaches such as Inceptiontime [7] have been proposed. Hybrid approaches like HIVE-COTE v2.0, combining multiple classifiers of different kind, are generally the most accurate, at a high computing cost [8]. For a recent general survey of TSC methods, the interested reader may refer to [4].
In the next section, we will focus on a specific type of TSC methods using random convolution kernels to extract features from time series data.
2.2 Random convolution kernels based methods
As mentioned above, these methods make use of random convolutions to discriminate time series. Convolution kernels can be seen as mini-random sequences whose convolution with time series generate activation maps used to extract several synthetic features that are subsequently fed into a simple linear classifier. The training phase thus only involves learning the appropriate weights assigned to each feature, which allows these methods to be, at the time this article is written, among the fastest while ensuring very good performance [4].
2.2.1 Main methods
ROCKET (Random Convolutional Kernel Transform) [1] is the first algorithm of this kind to have been introduced. It randomly generates a large number of convolution kernels (typically 10,000) which it uses to create activation maps. These maps are then summarized by two pooling operators: PPV (Proportion of Positive Values), which computes the percentage of positive values, and GMP (Global Maximum Pooling), which extracts the maximum value from the activation map. Formally:
where is the convolution output of .
For each time series, features are extracted, where is the number of kernels.
Although random, the kernels are parameterized as follows:
Length is randomly chosen from with equal probability.
Weights are randomly selected and then normalized, .
A bias is added to the activation map.
Dilations are computed with , where and , where is the length of the input time series, and the length of the kernel.
MiniRocket [2] is a variant of ROCKET with several key modifications that greatly speed-up the training without sacrificing performances. The method becomes much more deterministic by making the following changes:
Kernel length is fixed at 9 instead of .
It uses a single set of 84 kernels containing only as values. This major change optimizes convolution operations and significantly reduces computation time.
The bias is now derived from the convolution result with a kernel/dilation pair.
GMP is no longer calculated, only PPV is used.
MultiRocket [3] extends MiniRocket by introducing several improvements:
First-order difference () between two time units, corresponding to the rate of change, is used as an additional input representation. Formally:
(4) -
In addition to PPV, the following features are extracted: Mean of Positive Values (MPV), Mean of Indices of Positive Values (MIPV) and Longest Stretch of Positive Values (LSPV), formally defined by:
(5) where is the subset of positive value in ,
(6) where indicates the indices of positive values and
HYDRA (Hybrid Dictionary-ROCKET Architecture) [9] combines aspects of dictionary-based methods and random convolutions. The general idea is to group all kernels into groups of kernels. For each time series in the dataset, convolution is performed with all kernels from a given group. At each point in the series, the kernel leading to the highest convolution value is selected. Thus, for each group, a histogram of maximum kernel responses is created, serving as features for the classifier. First-order difference is used here as well. The kernels are parameterized as follows:
Length is fixed at 9.
Weights are randomly selected.
No bias or pooling operator is used after convolution.
The features generated by HYDRA and MultiRocket can be concatenated leading to a more accurate classification [9].
2.2.2 Kernel pruning
The main drawback of methods based on random convolutions is the large number of kernels, which may prevent them from being integrated into memory-limited devices. Moreover, not all kernels contribute positively to model performance, resulting in unnecessary computational complexity [10].
Various approaches have therefore been proposed for pruning the kernels, keeping only those most important to the problem.
S-ROCKET [10] selects the most important kernels and prunes redundant ones as follows:
1.
Pre-training the original model at full capacity,
2.
Initializing a population of binary candidate state vectors where each vector element represents the active/inactive status of a kernel,
3.
Using a population-based optimization algorithm to evolve this population, aiming to minimize the number of active kernels while maximizing classification accuracy,
4.
Combining the total number of active kernels with the classification accuracy to determine the best state vector, which is then used to select the optimal kernels.
Authors show that this approach maintains the classification performance of the original model while significantly reducing computational demands.
POCKET [11] improves S-ROCKET by dividing the pruning problem in 2 stages:
1.
Applying dynamically varying penalties for group-level regularization to prune redundant kernels.
2.
Uses element-level regularization to refit a linear classifier with the remaining features.
Experimental results demonstrate that POCKET performs 11 times faster than S-ROCKET while being as accurate.
Detach-ROCKET [12] prunes kernels using Sequential Feature Detachment (SFD), a method to identify and remove non-essential features. SFD involves iterative detachment steps, where a fixed percentage of the current active features is removed at each step. Depending on the number of classes, one or several RIDGE classifiers are trained on the set of active features, and RIDGE regressions are made to learn coefficients associated with each feature allowing one to eliminate the least relevant features.
2.2.3 ROCKET ensembles
Another way of improving a ROCKET classifier is to consider an ensemble of ROCKET methods.
Arsenal [8] is a set of small convolution-based ROCKET classifiers that produces a better probability estimate than ROCKET. It was introduced as a replacement of ROCKET in HIVE-COTE v2.0. As an ensemble method, Arsenal uses majority vote to classify new cases.
FT-FVC [13] adds feature diversity to ROCKET using three transformations to raw data: Hilbert and first-order and second-order difference. These transformations generate feature vectors that are concatenated with the raw feature vector to form a three-view representation. These representations are merged through hard voting.
3 On the importance of choosing the right input representations and pooling operator
To understand how input representations and pooling operators impact the performances of random convolution kernels based transforms, we compared, on the 112 datasets of the UCR archive [5], original MiniRocket with modified versions using GMP, MPV, MIPV and LSPV instead of PPV as pooling operator and the first order difference as additional input representation, leading to possible transforms (detailed in Section 4.1). In this article, transforms using multiple pooling operators were not considered since preliminary results showed they under-performed those using only one pooling operator in almost all the cases. Table 1 shows the mean classification accuracy across resamples of the possible transforms for some datasets included in UCR archive. For each dataset, the number in bold indicates the greatest overall accuracy achieved. Note that the PPV column is also the performance of the original MiniRocket classifier. Figure 1 shows the number of datasets in the UCR archive for which the modified MiniRocket version performs the best. If PPV_MIX is the best transform on average, it is outperformed by others in most cases ().
Figure 2 shows the critical difference diagram [14] of the 15 possible transforms for the different datasets, the distribution of classification accuracy is provided as box plots in Figure 13.
| DATASET | PPV | GMP | MPV | MIPV | LSPV | PPV | GMP | MPV | MIPV | LSPV | PPV | GMP | MPV | MIPV | LSPV |
| DIFF | DIFF | DIFF | DIFF | DIFF | MIX | MIX | MIX | MIX | MIX | ||||||
| ACSF1 | 82.77 | 55 | 82.97 | 69.23 | 85.27 | 79.9 | 60.6 | 81.8 | 68.6 | 84.4 | 82.2 | 61.4 | 83.03 | 70.47 | 85.57 |
| Adiac | 80.26 | 70.92 | 80.14 | 78.78 | 79.94 | 82.19 | 64.88 | 79.68 | 75.44 | 79.52 | 82.86 | 71.62 | 82.05 | 79.09 | 81.52 |
| ArrowHead | 88.15 | 86.34 | 87.6 | 85.85 | 88.91 | 87.9 | 83.41 | 85.77 | 81.22 | 88.57 | 88.7 | 86.67 | 88.06 | 85.05 | 89.49 |
| Beef | 77 | 61.44 | 75.22 | 73.56 | 68.67 | 77.78 | 66.22 | 71.56 | 77.78 | 67.56 | 79.67 | 68.89 | 74.56 | 77.44 | 70.56 |
| BeetleFly | 91 | 83 | 85.67 | 88 | 86.83 | 90.83 | 91.5 | 90.67 | 88.83 | 87.17 | 91.83 | 87.67 | 88.5 | 91 | 87.83 |
| BirdChicken | 92 | 90.33 | 89.83 | 83.67 | 91.33 | 90.67 | 93.33 | 91.83 | 88 | 89.83 | 90.17 | 91.33 | 88.5 | 85.33 | 90.17 |
| BME | 99.18 | 98.16 | 99.73 | 99.31 | 92.73 | 98.13 | 96.4 | 99.78 | 99.29 | 92.09 | 99.2 | 98.4 | 99.87 | 99.47 | 93.96 |
| Car | 91.78 | 87.28 | 91.44 | 91.61 | 88.83 | 91.06 | 76 | 89.17 | 88.89 | 83.94 | 92.61 | 86.33 | 92.33 | 92 | 89.33 |
| CBF | 99.62 | 98.7 | 99.4 | 96.09 | 99.39 | 83.23 | 76.34 | 80.93 | 79.71 | 81.63 | 99.36 | 97.69 | 99.24 | 94.38 | 98.88 |
| Chinatown | 96.87 | 95.17 | 96.66 | 97.28 | 96.14 | 95.83 | 90.39 | 94.88 | 96.36 | 94.7 | 96.56 | 94.19 | 96.03 | 97.14 | 95.96 |
| ChlorineConcent. | 75.42 | 73.45 | 71.75 | 75.86 | 73.29 | 78.69 | 82.53 | 78.07 | 78.55 | 75.64 | 78.15 | 81.86 | 77.59 | 78.65 | 75.81 |
| CinCECGTorso | 87.58 | 80.48 | 87.02 | 93.7 | 87.27 | 95.79 | 75.79 | 92.83 | 96.59 | 86.58 | 92.56 | 80.79 | 90.74 | 96.09 | 89.56 |
| Coffee | 99.88 | 99.88 | 99.88 | 100 | 100 | 98.69 | 97.26 | 99.88 | 99.64 | 98.57 | 99.76 | 99.29 | 100 | 100 | 100 |
| Computers | 80.16 | 73.12 | 81.44 | 72.89 | 79.36 | 84.84 | 67.35 | 82.23 | 74.96 | 79.95 | 84.87 | 73.87 | 84.56 | 75.75 | 82.6 |
| CricketX | 82.62 | 73.26 | 81.25 | 79.51 | 80.77 | 70.76 | 50.56 | 68.62 | 65.49 | 69.4 | 81.74 | 74.19 | 81.86 | 78.2 | 79.74 |
| CricketY | 84.24 | 74.61 | 82.51 | 79.46 | 81.45 | 68.76 | 45.69 | 67.48 | 65.8 | 68.08 | 83.51 | 75.06 | 82.92 | 80.08 | 80.78 |
| CricketZ | 84.25 | 75.86 | 83.42 | 81.2 | 82.44 | 72.18 | 49.69 | 70.23 | 65.42 | 71.68 | 83.14 | 76.36 | 83.69 | 79.66 | 82.17 |
| Crop | 76.37 | 69.89 | 76.17 | 75.94 | 76.25 | 70.65 | 63.26 | 71.04 | 70.66 | 70.25 | 76.71 | 71.86 | 76.99 | 76.12 | 76.37 |
| DiatomSizeReduction | 94.29 | 96.21 | 94.61 | 95.68 | 95.46 | 93.69 | 92 | 94.07 | 95.46 | 92.02 | 94.23 | 95.41 | 94.62 | 95.92 | 95.17 |
| DistalPha.Out.AgeGr. | 79.42 | 78.13 | 79.66 | 80.12 | 79.62 | 80.17 | 77.31 | 81.03 | 80.41 | 79.57 | 79.74 | 77.53 | 79.62 | 80.05 | 80.19 |
| DistalPha.Out.Corr. | 82.44 | 79.83 | 82.57 | 82.95 | 82.73 | 83.97 | 80.92 | 83.32 | 83.85 | 82.8 | 83.93 | 81.44 | 83.67 | 83.8 | 83.47 |
| DistalPhalanxTW | 69.5 | 67.84 | 69.81 | 69.47 | 69.06 | 69.9 | 67.39 | 69.95 | 68.25 | 68.9 | 69.81 | 67.63 | 69.66 | 68.85 | 68.94 |
| Earthquakes | 73.79 | 73.45 | 73.93 | 74.96 | 73.96 | 74.6 | 71.7 | 74.2 | 74.82 | 74.82 | 74.12 | 73.74 | 74.15 | 75.2 | 74.44 |
| ECG200 | 90.2 | 88.57 | 89.67 | 88.43 | 89.17 | 84.63 | 83.17 | 85.07 | 83.9 | 82.43 | 89.67 | 86.07 | 89.17 | 87.6 | 87.77 |
| ECG5000 | 94.65 | 94.64 | 94.74 | 94.56 | 94.53 | 94.21 | 94.03 | 94.41 | 94.26 | 94.18 | 94.65 | 94.64 | 94.74 | 94.57 | 94.6 |
| ECGFiveDays | 99.07 | 99.69 | 99.65 | 97.5 | 99.55 | 98.83 | 95.74 | 99.04 | 96.88 | 98.35 | 99.21 | 99.77 | 99.73 | 97.23 | 99.66 |
| ElectricDevices | 87.44 | 74.46 | 87.71 | 85.12 | 87.07 | 85.83 | 69.01 | 85.47 | 79.61 | 84.79 | 89.16 | 76.26 | 88.99 | 86 | 88.13 |
| EOGHorizontalSignal | 83.46 | 79.91 | 84.66 | 80.9 | 83.49 | 80.02 | 65.91 | 80.71 | 73.49 | 76.73 | 85.17 | 78.8 | 86.56 | 81.57 | 84.79 |
| EOGVerticalSignal | 79.91 | 76.02 | 80.9 | 76.77 | 79.24 | 73.78 | 58.79 | 75.78 | 68.69 | 73.27 | 79.4 | 74.19 | 82.14 | 76.57 | 79.25 |
| EthanolLevel | 66.27 | 58.43 | 62.74 | 71.87 | 70.86 | 60.56 | 37.51 | 51.85 | 65.53 | 51.55 | 66.75 | 56.09 | 62.42 | 71.49 | 69.89 |
| FaceAll | 98.35 | 96.96 | 98.65 | 98.31 | 97.96 | 96.88 | 93.2 | 97.07 | 97.11 | 96 | 98.36 | 96.91 | 98.64 | 98.26 | 97.81 |
| FaceFour | 94.24 | 81.25 | 92.31 | 90.08 | 92.99 | 78.03 | 59.39 | 75.3 | 81.14 | 78.56 | 92.8 | 77.58 | 89.62 | 88.64 | 90.38 |
| FacesUCR | 96.92 | 93.5 | 96.71 | 97.11 | 95.87 | 92.4 | 86.88 | 93.13 | 93.54 | 90.93 | 96.6 | 93.57 | 96.63 | 96.77 | 95.2 |
| FiftyWords | 82.89 | 72.59 | 81.53 | 82.32 | 79.71 | 80.13 | 53.93 | 74.44 | 79.42 | 74.63 | 83.69 | 72.64 | 81.25 | 82.31 | 79.68 |
4 SelF-Rocket
Following the experimental results obtained in the previous section, we investigate a new variant of ROCKET, called SelF-Rocket (Selected Features Rocket), that aims at dynamically selecting the best couple of input representations and pooling operator during the training process. The latter involves three main stages:
SelF-Rocket works in a similar way to MiniRocket or MultiRocket. The only major difference between them is the addition of a feature selection phase. As seen previously in Section 3, the optimal set of features varies from one dataset to another, so using a single feature type (PPV) like MiniRocket or a fixed set of feature types ({PPV, LSPV, MIPV, MPV}) like MultiRocket can produce good results on average, but not optimal for each case. To address this problem, we implement a wrapper-based feature selection module with the objective to select the most appropriate set for each classification problem.
4.1 Feature Generation
Let IR be a set of Input Representations, PO a set of Pooling Operators, MK the set of all kernels (containing every kernel/dilation combinations) generated by MiniRocket and the bias associated to a kernel (cf Section 2.2.1).
In the implementation evaluated in this article we consider
and
where is the identity function ( for all time series ) and is the first order difference (cf Equation 4).
Let be a feature generation function defined as:
where is a time series, a subset of the power set of IR (minus the empty set), i.e. , and a pooling operator, i.e. .
Thus, following this formalization, the original MiniRocket extracts features using the parametrization , for any time series .
The Cartesian product induces a set of such parameterized functions; In this implementation, and , leading to possible parametrizations.
Let FV be the generated Feature Vector. In the following, feature generation outputs are denoted as in Section 3 for readability reasons:
p for ,
p_DIFF for ,
and p_MIX for ,
for all .
4.2 Feature Selection
The Feature Selection Module relies on the Stratified -Fold method to create novel train and validation sets derived from the original train set. These sets are then used to train mini-classifiers for each IR-PO combination. Subsequently, the optimal combination on average is selected as the final set of features for the linear classifier. The Feature Selection Module is summarized in Algorithm 1.
Training set class y_train
The Number of folds k
The Number of features per mini-classifier f
4.3 Classification
SelF-Rocket employs the same algorithm for the mini-classifiers embedded within the Feature Selection Module and the end-stage classifier, namely the RIDGE classifier, as proposed by Dempster et al. [1]. This classifier is preferable to stochastic descent methods, such as logistic regression, when the number of features exceeds the number of training examples and when working with small datasets, e.g. when there are fewer than 10,000 examples.
5 Experiments
In this section, the performances of SelF-Rocket are evaluated.
We show that SelF-Rocket is as accurate as Hydra-MultiRocket and HIVE-COTE v2.0, despite having only one type of feature, and also has a relatively shorter classifier training time. We also investigate the
influence of the main parameters on the model performance.
5.1 Experimental settings
SelF-Rocket has been evaluated on a subset of the UCR Time Series Archive [5], which is widely used in this field. To guarantee the reproducibility of our experiments and a fair comparison with other TSC algorithms, we employed the identical 112 datasets as those used by [4] and the precise same 30 train/test resamples for each of those datasets. The 112 datasets are those from the 128 original ones that did not contain missing values or variable time series length.
Note that we did not compute the results for others methods shown in Figures 4, 5 and 6, but used the results111https://github.com/time-series-machine-learning/tsml-eval/tree/main/results/classification/Univariate available in [4].
The author’s implementation of MiniRocket222https://github.com/angus924/minirocket/tree/main was used as baseline model. Our algorithm is implemented in standard python 3.11.4.
In order to compare the classification performance with other TSC algorithms, we use the critical difference diagrams that display mean ranks of each method across all datasets with the horizontal cliques indicating that there is no statistically significant difference between those methods, the Multiple Comparison Matrix (MCM) [15] with heat-map color representing mean differences in score, and the pairwise scatter plots of test accuracy summarizing the number of win/draw/loss between 2 classifiers across all datasets.
All experiments were conducted on a Dell Precision 5480 on Debian 12 with an Intel Core i9-13900H vPro 2.6 GHz and 32GB of RAM
5.2 Comparison with other TSC methods
We compare the performances of SelF-Rocket with 11 TSC methods: 1NN-DTW, Elastic Ensemble (EE) [16], Shapelet transform classifier (STC) [17], HIVE-COTE v2.0 [8], H-Inceptiontime [18], MultiRocket [3], Hydra, Hydra+MultiRocket [9], MiniRocket [2],
ROCKET [1] and QUANT [19].
Figure 4 shows the average mean accuracy rank of these classifiers. In terms of performance, SelF-Rocket ranks second only to HIVE-COTE v2.0, and outperforms others.
Figure 5 shows the Multiple Comparison Matrix [20], i.e. the full pairwise comparison between the best performing methods tested. The mean difference in accuracy between SelF-Rocket and its baseline MiniRocket is approximately 1.23 %
Figure 6 displays the mean accuracy, the pairwise win/draw/loss, and value for statistical tests between the tested methods over 30 resamples for the 112 selected UCR datasets. Each point represents a dataset, the more a point is distant to the diagonal line, the more one of the two methods performs better than the other. In some cases, SelF-Rocket outperforms Hydra-MultiRocket, but in general, both methods offer comparable performance. HIVE-COTE v2.0 also has the advantage over SelF-Rocket in certain datasets.
There is no significant statistical difference by Wilcoxon signed rank test between SelF-Rocket, Hydra-MultiRocket and HIVE-COTE v2.0
SelF-Rocket (9,996 or 19,992 features) is faster in the training and prediction phase than MultiRocket and Hydra-MultiRocket due to a lower number of features (resp 50,000 and 50,000 + number of features of Hydra).
5.3 Sensitivity Analysis and Ablative Study
As previously shown, SelF-Rocket outperforms other existing methods based on random convolution kernels. In this section, the impact of the three key parameters on global accuracy performance is studied:
Following [21, 19], we choose 50 ‘development’ datasets from the 112 UCR datasets to conduct this analysis. We first remove the 12 datasets with the most training examples, then we select the sets that contain at least 5 examples per class and 100 training examples. From the remaining list, 50 were randomly selected. The complete list is available in Table 4.
The others parameters of SelF-Rocket, i.e. number of kernels, padding, dilation, bias, length of kernel, values inside the kernel, remain the same as the default parameters of MiniRocket.
5.3.1 Number of folds
The SelF-Rocket Feature Selection module employs a stratified -fold decomposition of the training set. We have conducted tests using the minimal number of folds possible, namely , up to a fixed value of . Figure 7 illustrates the impact of the number of folds on the classifier mean accuracy. No notable differences were observed when different values of were used.
5.3.2 Number of features
In its default configuration, MiniRocket only generates 9,996 kernels using PPV as pooling operator, whereas SelF-Rocket employs 9,996 kernels with the base input representation and an additional 9,996 with the first-order difference input representation.
For each couple (input representation, pooling operator) belonging to , SelF-Rocket generates 9,996 features. The number of features tested ranges from 2,500 to 9,996 for each mini-classifier within the Feature Selection module, features are randomly selected from the total number of features, which is either 9,996 or 19,992 ().
Figure 7 shows that the number of features has a relatively minor impact on the average accuracy of the final classifier.
5.3.3 Input Representations and Pooling Operators
| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
| ONLY BASE | ONLY DIFF | ONLY MIX | ||
| Win | – | 2 | 4 | 6 |
| Draw | – | 18 | 15 | 23 |
| Lose | – | 30 | 31 | 21 |
| Mean accuracy | 85.69 | 85.01 | 82.13 | 85.35 |
We evaluated the classification performance of the features generated by 3 sets of input representations: the raw time series (}), the first-order difference applied to the time series (}), and a concatenation of features for both (}). Figure 8 illustrates the discrepancy in ranking between the utilization of all three sets of representations and the exclusive use of a single one.
Table 2 provides a summary of the pairwise win/draw/loss between SelF-Rocket () and SelF-Rocket with only one set of representations.
The first-order difference is generally the least impactful representation, while the MIX set of representations demonstrates the best overall performance among the three.
| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
| NO PPV | NO GMP | NO MIPV | NO MPV | NO LSPV | ||
| Win | – | 4 | 1 | 1 | 0 | 4 |
| Draw | – | 27 | 48 | 39 | 46 | 43 |
| Lose | – | 19 | 1 | 10 | 4 | 3 |
| Mean acc. | 85.69 | 85.44 | 85.63 | 85.36 | 85.60 | 85.67 |
The choice of a pooling operator or a set of pooling operators to generate features can affect the performance of the classifier in the classification process. Figure 9 and Table 3 show the impact on performance of not using one of the five pooling operators. It can be seen that not using GMP or LSPV has no significant impact on system performance. On the other hand, not using MIPV or PPV has a greater impact on overall performance.
5.4 Scalability and Compute Time Comparison
Compared to other ROCKET methods, SelF-Rocket incorporates a Feature Selection module between the global set of features and the final classifier. The time required for the generation of SelF-Rocket features is less than Hydra-MultiRocket, while the training of the classifier is similar in duration to that of MiniRocket.
Figure 10 shows the mean compute time for all resamples and all datasets of the Feature Selection module according to the number of folds and features. The classifier complexity depends on , the number of features and , the number of training examples. is a multiplier for the number of mini-classifiers, and impacts the complexity of a single classifier so the two parameters impact generally the compute time.
For each combination of IR and PO, we instantiate mini-classifiers to identify the optimal set of features. A RIDGE classifier with a complexity of is employed. In Figure 11(a) given that is a fixed value, namely 9,996, we can see that the mean time only depends on the number of training examples.
Figure 12(a) shows the average total computation time over the 50 development datasets, for all resamples. SelF-Rocket and MiniRocket training (Figure 12(a)) and prediction times (Figure 12(b)) are comparable. However, among the evaluated methods, MiniRocket remains the fastest overall. As illustrated in Figure 10, with and SelF-Rocket’s average feature selection time is around 1.3 seconds, which is significantly faster than its feature creation computation time. SelF-Rocket features take longer to create than MultiRocket features. This is due to the greater number of kernels that are instantiated. MultiRocket instantiates only 6,216 kernels for each of its input representations (DIFF and BASE). From these, features () are generated. In comparison, SelF-Rocket generates 9,996 kernels for each of these two representation types for a total of features ().
SelF-Rocket produces times more features than MiniRocket. One way to improve its overall compute time could be to reduce the number of kernels generated by SelF-Rocket or to reduce the number of input representations and pooling operators.
HYDRA produces kernels : (number of kernels per group), (number of groups) and (maximum possible dilation value), with 2 features extracted per kernel (maximum and minimum responses). For each dataset, 512 ( and ) kernels per dilation are created with . In most of our cases will be between 7 and 10, so the total number of features generated will generally be between (7,128) and (10,240). This explains why HYDRA is the fastest in terms of classifier training and testing time. As illustrated in Figures 12(c) and 12(d), SelF-Rocket is faster than Hydra-MultiRocket in terms of both feature creation time and classifier testing time, with less discrepancy.
6 Conclusion and prospects
The two main contributions of this article are as follows:
First, we have shown that the choice of Input Representations and Pooling Operators significantly impacts the performance of TSC methods based on random convolution kernels. We have also shown that using multiple Pooling Operators, such as MultiRocket, leads to good performances on average, but is not optimal in specific cases: it is always preferable to select a single appropriate Pooling Operator.
Then, on this basis, we have proposed a new algorithm based on MiniRocket, SelF-Rocket, incorporating a feature selection stage. Experiments show that this algorithm selects the best Input Representations / Pooling Operator combination in most cases, leading to state of the art performances. In the tested implementation of SelF-Rocket, only 9,996 or 19,992 features are retained to train the RIDGE classifier, far fewer than competing methods such as MultiRocket or Hydra-MultiRocket, leading to fast predictions.
In this article, a minimal version of SelF-Rocket was evaluated, using 2 input representations, 5 pooling operators and a simple wrapper-based feature selection method. This implementation could be improved by:
Considering more input representations (second order difference, Fourier and Hilbert transforms, etc.),
Combining SelF-Rocket and HYDRA features,
Using HYDRA as a pooling operator within SelF-Rocket,
Using a filter-based feature selection method (based on test for instance) instead of Stratified -Fold to speed-up the feature selection stage.
Acknowledgments
This research project, supported and financed by the French ANR (Agence Nationale pour la Recherche), is part of the Labcom (Laboratoire Commun) MYEL (MobilitY and Reliability of Electrical chain Lab) involving LSEE, LGI2A and CRITTM2A (ANR-22-LCV2-0001 MYEL).
Appendix A Additional Figures & Tables
| Developement datasets | ||
| ACSF1 | ChlorineConcentration | Computers |
| CricketX | CricketY | CricketZ |
| DistalPhalanxOutlineAgeGroup | DistalPhalanxOutlineCorrect | DistalPhalanxTW |
| Earthquakes | ECG200 | EOGHorizontalSignal |
| EOGVerticalSignal | EthanolLevel | FaceAll |
| Fish | FreezerRegularTrain | GunPointAgeSpan |
| GunPointMaleVersusFemale | GunPointOldVersusYoung | Ham |
| Haptics | InlineSkate | InsectWingbeatSound |
| LargeKitchenAppliances | MedicalImages | MiddlePhalanxOutlineAgeGroup |
| MiddlePhalanxOutlineCorrect | MixedShapesRegularTrain | MixedShapesSmallTrain |
| OSULeaf | Plane | PowerCons |
| ProximalPhalanxOutlineAgeGroup | ProximalPhalanxOutlineCorrect | ProximalPhalanxTW |
| ScreenType | SemgHandGenderCh2 | SemgHandMovementCh2 |
| SemgHandSubjectCh2 | ShapesAll | SmoothSubspace |
| SwedishLeaf | SyntheticControl | Trace |
| UWaveGestureLibraryX | UWaveGestureLibraryY | UWaveGestureLibraryZ |
| Worms | WormsTwoClass | |

### v4

@@
 Artois, UR 4025, Laboratoire Systèmes Electrotechniques et Environnement (LSEE), F-62400 Béthune, France 2]Univ.
-Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels based transforms: pooling operators and input representations matter
+Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels: pooling operators and input representations matter
 Abstract
-This article presents a new approach based on MiniRocket, called SelF-Rocket, for fast time series classification (TSC).
-Unlike existing approaches based on random convolution kernels, it dynamically selects the best couple of input representations and pooling operator during the training process.
+This article presents a new approach based on MiniRocket, called SelF-Rocket (Selected Features Rocket), for fast time series classification (TSC).
+Unlike existing approaches based on random convolution kernels, it dynamically attempts to select the best couple of input representations and pooling operator during the training process.
 SelF-Rocket achieves state-of-the-art accuracy on the University of California Riverside (UCR) TSC benchmark datasets.
@@
 Time series classification, ROCKET, MiniRocket, Pooling operators, Input representation, Feature selection.1 Introduction
-Random convolution kernels based transforms, ROCKET [1] and its successors, MiniRocket [2], MultiRocket [3] and others, have revolutionized the field of time series classification, offering an unrivaled compromise between speed and accuracy [4].
+Random convolution kernels based transforms, ROCKET [1] and its successors, MiniRocket [2], MultiRocket [3] and others, have significantly contributed to the advancement of time series classification, offering an unrivaled compromise between speed and accuracy [4].
 The main idea behind ROCKET methods is to use a very large number of random convolution kernels to transform the time series and apply one or more pooling operators to generate a set of features.
@@
 To the best of our knowledge, all the proposed transformations rely on one or more fixed pooling operators such as, for example, PPV (Proportion of positive values), for feature extraction from time series.
-In this article, a new approach based on MiniRocket, called SelF-Rocket, is developed.
-Unlike existing ones based on random convolution kernels, it dynamically selects the best data input representations and pooling operator combination during the training process.
+In this article, a new approach based on MiniRocket, called SelF-Rocket (Selected Features Rocket), is developed.
+Unlike existing ones based on random convolution kernels, it dynamically attempts to select the best couple (or combination) of data input representations and pooling operator during the training process.
+The complete source code of our implementation of SelF-Rocket is available on GitHub111https://github.com/ANR-MYEL/SelF-Rocket/.
 This article is organized as follows.
-The fundamentals of time series classification and the main random convolution kernels based methods are recalled in Section 2.
+The fundamentals of time series classification and the main random convolution kernels based methods are first recalled in Section 2.
 A statistical analysis, conducted on the University of California Riverside (UCR) archive [5], is then presented in Section 3 to motivate this work.
@@
 The method SelF-Rocket is then introduced in Section 4 and its performances on the UCR archive are analyzed in Section 5.
-Finally, Section 6 offers a conclusion with some perspectives.
-2 Related works
-2.1 Time series classification
+Finally, Section 6 gives a conclusion with some perspectives.
+2 Time series classification
+2.1 Overview
 Time Series Classification (TSC) is a type of supervised learning problem where the objective is to assign a class label to a given time series instance.
-A Time series data consist of sequences of observations collected at constant intervals over time.
+The data in a time series is made up of sequences of observations collected at constant intervals over time.
 Formally, a time series can be represented as:
 where denotes the value of the time series at time step , and represents the length of the time series. This notation will be used to refer to a time series throughout this article.
-Many methods of various kind have been developed to address TSC problems.
+Many methods of various kinds have been developed to address TSC problems.
 Traditional methods include distance-based techniques like Dynamic Time Warping (DTW) [6], which measure the similarity between time series.
 More recently, Deep Learning approaches such as Inceptiontime [7] have been proposed.
-Hybrid approaches like HIVE-COTE v2.0, combining multiple classifiers of different kind, are generally the most accurate, at a high computing cost [8].
+Hybrid approaches like HIVE-COTE v2.0, combining multiple classifiers of different kinds, are generally the most accurate, at a high computing cost [8].
 For a recent general survey of TSC methods, the interested reader may refer to [4].
-In the next section, we will focus on a specific type of TSC methods using random convolution kernels to extract features from time series data.
+The next section focuses on a particular type of CST method that involves the use of random convolution kernels for feature extraction from time series data.
 2.2 Random convolution kernels based methods
@@
 where is the convolution output of .
-For each time series, features are extracted, where is the number of kernels.
+For each kernel, two features are extracted using PPV and GMP.
 Although random, the kernels are parameterized as follows:
 Length is randomly chosen from with equal probability.
-Weights are randomly selected and then normalized, .
+Weights are randomly selected and then normalized: .
 A bias is added to the activation map.
@@
 (5) where is the subset of positive value in ,
-(6) where indicates the indices of positive values and
+(6) where indicates the indices of positive values, and
 HYDRA (Hybrid Dictionary-ROCKET Architecture) [9] combines aspects of dictionary-based methods and random convolutions.
@@
 2.2.2 Kernel pruning
-The main drawback of methods based on random convolutions is the large number of kernels, which may prevent them from being integrated into memory-limited devices.
+The primary limitation of methods that rely on random convolutions is the extensive number of kernels they require, which may prevent them from being integrated into memory-limited devices.
 Moreover, not all kernels contribute positively to model performance, resulting in unnecessary computational complexity [10].
-Various approaches have therefore been proposed for pruning the kernels, keeping only those most important to the problem.
+As a result, several methods have been suggested for pruning the kernels, retaining only those that are most relevant to the problem.
 S-ROCKET [10] selects the most important kernels and prunes redundant ones as follows:
 1.
-Pre-training the original model at full capacity,
+The original model is initially pre-trained to its full capacity.
 2.
-Initializing a population of binary candidate state vectors where each vector element represents the active/inactive status of a kernel,
+A population of binary candidate state vectors is initialized, where each element in a vector indicates whether a kernel is active or inactive.
 3.
-Using a population-based optimization algorithm to evolve this population, aiming to minimize the number of active kernels while maximizing classification accuracy,
+A population-based optimization algorithm is employed to evolve this population, with the goal of reducing the number of active kernels while enhancing classification accuracy.
 4.
-Combining the total number of active kernels with the classification accuracy to determine the best state vector, which is then used to select the optimal kernels.
+The total number of active kernels and the classification accuracy are combined to identify the best state vector, which is then used to select the optimal set of kernels.
 Authors show that this approach maintains the classification performance of the original model while significantly reducing computational demands.
-POCKET [11] improves S-ROCKET by dividing the pruning problem in 2 stages:
+POCKET [11] enhances S-ROCKET by addressing the pruning problem in two distinct stages:
 1.
-Applying dynamically varying penalties for group-level regularization to prune redundant kernels.
+It applies dynamically varying penalties for group-level regularization to remove redundant kernels.
 2.
-Uses element-level regularization to refit a linear classifier with the remaining features.
-Experimental results demonstrate that POCKET performs 11 times faster than S-ROCKET while being as accurate.
-Detach-ROCKET [12] prunes kernels using Sequential Feature Detachment (SFD), a method to identify and remove non-essential features.
+It employs element-level regularization to refit a linear classifier using the remaining features.
+Experimental results show that POCKET performs 11 times faster than S-ROCKET while being as accurate.
+Detach-ROCKET [12] prunes kernels using Sequential Feature Detachment (SFD), a technique to identify and remove non-essential features.
 SFD involves iterative detachment steps, where a fixed percentage of the current active features is removed at each step.
@@
 As an ensemble method, Arsenal uses majority vote to classify new cases.
-FT-FVC [13] adds feature diversity to ROCKET using three transformations to raw data: Hilbert and first-order and second-order difference.
+FT-FVC [13] adds feature diversity to ROCKET using three transformations to raw data: Hilbert, first-order and second-order difference.
 These transformations generate feature vectors that are concatenated with the raw feature vector to form a three-view representation.
-These representations are merged through hard voting.
+Three classifiers are trained separately on each view, and a prediction is made by hard voting.
 3 On the importance of choosing the right input representations and pooling operator
-To understand how input representations and pooling operators impact the performances of random convolution kernels based transforms, we compared, on the 112 datasets of the UCR archive [5], original MiniRocket with modified versions using GMP, MPV, MIPV and LSPV instead of PPV as pooling operator and the first order difference as additional input representation, leading to possible transforms (detailed in Section 4.1).
-In this article, transforms using multiple pooling operators were not considered since preliminary results showed they under-performed those using only one pooling operator in almost all the cases.
-Table 1 shows the mean classification accuracy across resamples of the possible transforms for some datasets included in UCR archive.
-For each dataset, the number in bold indicates the greatest overall accuracy achieved.
+To assess the influence of input representations and pooling operators on the performance of transformations based on random convolution kernels, we conducted a comparison using 112 selected datasets of the UCR archive [5].
+This comparison involved the original MiniRocket and modified versions, which utilized GMP, MPV, MIPV, and LSPV as pooling operators instead of PPV.
+Additionally, we incorporated the first-order difference (DIFF) as an extra input representation.
+Classifiers are denoted using a notation PO_IR depending on the pooling operator PO used () and the input representation IR used (MIX denoting the concatenation of the first order difference DIFF and the base representation).
+We then compared possible transformations, which are formally defined in Section 4.1.
+These transformations correspond to combinations of these 5 potential pooling operators with these 3 possible input representations: the base representation, Diff, and MIX.
+It is worth mentioning that we also conducted experiments using multiple pooling operators, rather than selecting just one as described here.
+However, our preliminary results suggested that using multiple operators did not yield performances as effective as those discussed in this section.
+Table 1 illustrates the mean classification accuracy across resamples of some datasets included in UCR archive for the possible transforms.
+For each dataset, the number in bold indicates the highest overall accuracy achieved.
 Note that the PPV column is also the performance of the original MiniRocket classifier.
@@
 If PPV_MIX is the best transform on average, it is outperformed by others in most cases ().
-Figure 2 shows the critical difference diagram [14] of the 15 possible transforms for the different datasets, the distribution of classification accuracy is provided as box plots in Figure 13.
+Figure 2 displays the critical difference diagram [14] of the 15 possible transforms for the different datasets.
+In Appendix A, the distribution of classification accuracy is also provided as box plots in Figure 18.
 | DATASET | PPV | GMP | MPV | MIPV | LSPV | PPV | GMP | MPV | MIPV | LSPV | PPV | GMP | MPV | MIPV | LSPV |
@@
 | FiftyWords | 82.89 | 72.59 | 81.53 | 82.32 | 79.71 | 80.13 | 53.93 | 74.44 | 79.42 | 74.63 | 83.69 | 72.64 | 81.25 | 82.31 | 79.68 |
+Based on these experiments, it is evident that the combination of input representations and pooling operators yielding the best performance varies depending on the dataset.
+Therefore, it would be beneficial to have the capability to select the appropriate combination based on a training dataset.
+This is the goal of the SelF-Rocket method, which will be presented in the following section.
 4 SelF-Rocket
-Following the experimental results obtained in the previous section, we investigate a new variant of ROCKET, called SelF-Rocket (Selected Features Rocket), that aims at dynamically selecting the best couple of input representations and pooling operator during the training process.
-The latter involves three main stages:
-SelF-Rocket works in a similar way to MiniRocket or MultiRocket.
-The only major difference between them is the addition of a feature selection phase.
-As seen previously in Section 3, the optimal set of features varies from one dataset to another, so using a single feature type (PPV) like MiniRocket or a fixed set of feature types ({PPV, LSPV, MIPV, MPV}) like MultiRocket can produce good results on average, but not optimal for each case.
-To address this problem, we implement a wrapper-based feature selection module with the objective to select the most appropriate set for each classification problem.
+Following the experimental results obtained in the previous section, we investigate a new variant of ROCKET, called SelF-Rocket (Selected Features Rocket), that aims to dynamically select the best couple of input representations and pooling operator during the training process.
+The latter, illustrated in Figure 3, involves three main stages:
+SelF-Rocket operates similarly to MiniRocket or MultiRocket, the main difference being the inclusion of a feature set selection step.
+As previously observed in Section 3, the optimal set of features can differ from one dataset to another.
+Therefore, while using a single feature type like PPV, as in MiniRocket, or a fixed set of feature types such as {PPV, LSPV, MIPV, MPV}, as in MultiRocket, can yield generally good results, it may not be optimal for every specific case.
+To tackle this problem, we have incorporated a wrapper-based Feature Selection Module with the objective to select the most appropriate set of features for each classification problem.
 4.1 Feature Generation
-Let IR be a set of Input Representations, PO a set of Pooling Operators, MK the set of all kernels (containing every kernel/dilation combinations) generated by MiniRocket and the bias associated to a kernel (cf Section 2.2.1).
+Let IR be a set of Input Representations, PO a set of Pooling Operators, MK the set of all kernels (containing every kernel/dilation combination) generated by MiniRocket and the bias associated to a kernel (cf Section 2.2.1).
 In the implementation evaluated in this article we consider
@@
 where is the identity function ( for all time series ) and is the first order difference (cf Equation 4).
-Let be a feature generation function defined as:
+Let be a feature set generation function defined as:
 where is a time series, a subset of the power set of IR (minus the empty set), i.e. , and a pooling operator, i.e. .
@@
 4.2 Feature Selection
-The Feature Selection Module relies on the Stratified -Fold method to create novel train and validation sets derived from the original train set.
+Summarized in Figure 4 and Algorithm 1, the Feature Selection Module relies on stratified train/test split methods to generate new train and validation sets derived from the original train set.
+Depending on the size of that dataset, we use either a stratified -fold repeated times or a stratified shuffle split with splits.
+A stratified shuffle split is used if the dataset contains at least a sufficiently large number of data (denoted by in Algorithm 1), it should be noted that this value has a direct impact on the time available for finding the best IR-PO combination by limiting the maximum size of each train and validation set, and thus the time available for this task.
+This strategy ensures that, as the size of the dataset increases, the learning and validation sets will vary more significantly from one voter to another.
+It therefore allows more reliable assessments of the performance of the different combinations.
 These sets are then used to train mini-classifiers for each IR-PO combination.
-Subsequently, the optimal combination on average is selected as the final set of features for the linear classifier.
-The Feature Selection Module is summarized in Algorithm 1.
-Training set class y_train
-The Number of folds k
-The Number of features per mini-classifier f
+Subsequently, the selection of the optimal IR-PO combination across all the original train set splits is made using a highest median voting system, which is validated by the Algorithm 2 prior to its use as the final set of features for the linear classifier.
+Using the highest median to select an IR-PO as an alternative to the highest mean is preferable as it is more robust to extreme values (e.g.
+lucky runs) that may occur.
+The idea of the Algorithm 2 is to check if the selected IR-PO combination is sufficiently supported by the voters, otherwise a default combination is chosen.
+Its aim is to avoid choosing an IR-PO combination that has poor generalisability, especially with small datasets.
+In our implementation, Algorithm 2 checks if the selected IR-PO combination is part of the top values for each voter (e.g.
+in the top for each voter) with a certain degree of flexibility (e.g.
+at , implemented as threshold in Algorithm 2), otherwise the selected IR-PO combination is replaced by a default one (PPV_MIX in our case).
+Training set class
+The number of features per mini-classifier
+The number of runs
+The maximum dataset size
+Index of the chosen IR-PO combination
+The number of voters
+The top considered
+The value of the threshold
 4.3 Classification
@@
 5 Experiments
-In this section, the performances of SelF-Rocket are evaluated.
-We show that SelF-Rocket is as accurate as Hydra-MultiRocket and HIVE-COTE v2.0, despite having only one type of feature, and also has a relatively shorter classifier training time.
-We also investigate the
-influence of the main parameters on the model performance.
+In this section, the performances of SelF-Rocket and Hydra + SelF-Rocket (concatenation of the features of Hydra and SelF-Rocket) are evaluated.
+We show that SelF-Rocket is as accurate as MultiRocket, despite the use of only one pooling operator at a time, and also has a relatively shorter classifier prediction time.
+We also investigate the influence of the main parameters of SelF-Rocket on its performance.
 5.1 Experimental settings
-SelF-Rocket has been evaluated on a subset of the UCR Time Series Archive [5], which is widely used in this field.
-To guarantee the reproducibility of our experiments and a fair comparison with other TSC algorithms, we employed the identical 112 datasets as those used by [4] and the precise same 30 train/test resamples for each of those datasets.
+To guarantee the reproducibility of our experiments and a fair comparison with other TSC algorithms, the proposed method is evaluated on the same same identical 112 datasets of the UCR Time Series Archive [5] and the precise same 30 train/test resamples for each of those datasets [1, 9, 15, 3, 16, 4].
 The 112 datasets are those from the 128 original ones that did not contain missing values or variable time series length.
-Note that we did not compute the results for others methods shown in Figures 4, 5 and 6, but used the results111https://github.com/time-series-machine-learning/tsml-eval/tree/main/results/classification/Univariate available in [4].
-The author’s implementation of MiniRocket222https://github.com/angus924/minirocket/tree/main was used as baseline model. Our algorithm is implemented in standard python 3.11.4.
-In order to compare the classification performance with other TSC algorithms, we use the critical difference diagrams that display mean ranks of each method across all datasets with the horizontal cliques indicating that there is no statistically significant difference between those methods, the Multiple Comparison Matrix (MCM) [15] with heat-map color representing mean differences in score, and the pairwise scatter plots of test accuracy summarizing the number of win/draw/loss between 2 classifiers across all datasets.
-All experiments were conducted on a Dell Precision 5480 on Debian 12 with an Intel Core i9-13900H vPro 2.6 GHz and 32GB of RAM
+The evaluation of a TSC method on UCR may be limited [5, 17] (for example, the data are pre-processed and of the same length), nevertheless, as described above, it currently remains the standard for benchmarking TSC methods.
+Note that we did not compute the results for other methods shown in Figures 5, 6 and 7, but used the results222https://github.com/time-series-machine-learning/tsml-eval/tree/main/results/classification/Univariate available in [4].
+The original implementations of MiniRocket333https://github.com/angus924/minirocket and MultiRocket444https://github.com/ChangWeiTan/MultiRocket were used as baseline.
+Our algorithm is implemented in standard Python 3.11.4.
+In order to compare the classification performance with other TSC algorithms, we use the critical difference diagrams that display mean ranks of each method across all datasets with the horizontal cliques indicating that there is no statistically significant difference between those methods, the Multiple Comparison Matrix (MCM) [18] with heatmap color representing mean differences in score, and the pairwise scatter plots of test accuracy summarizing the number of win/draw/loss between 2 classifiers across all datasets.
+All experiments were conducted on a cluster using Ubuntu 24.04.1 LTS with an Intel(R) Xeon(R) Gold 6434 and 250GB of RAM.
+A sensitivity analysis of the main parameters of Self-Rocket is given in Section 5.3.
+We have chosen as default parameters for SelF-Rocket, a number of folds of , a number of features per mini-classifier of , a number of runs of , and a validation vote with a top and a threshold of .
+The others parameters of SelF-Rocket, i.e.
+number of kernels, padding, dilation, bias, length of kernel, values inside the kernel, remain the same as the default parameters of MiniRocket.
 5.2 Comparison with other TSC methods
-We compare the performances of SelF-Rocket with 11 TSC methods: 1NN-DTW, Elastic Ensemble (EE) [16], Shapelet transform classifier (STC) [17], HIVE-COTE v2.0 [8], H-Inceptiontime [18], MultiRocket [3], Hydra, Hydra+MultiRocket [9], MiniRocket [2],
+We compare the performances of SelF-Rocket and Hydra + SelF-Rocket with 11 TSC methods: 1NN-DTW, Elastic Ensemble (EE) [19], Shapelet transform classifier (STC) [20], HIVE-COTE v2.0 [8], H-Inceptiontime [21], MultiRocket [3], Hydra, Hydra+MultiRocket [9], MiniRocket [2],
 ROCKET [1] and QUANT [15].
-Figure 4 shows the average mean accuracy rank of these classifiers. In terms of performance, SelF-Rocket ranks second only to HIVE-COTE v2.0, and outperforms others.
-Figure 5 shows the Multiple Comparison Matrix [20], i.e.
+| Mean-Accuracy | HC2 0.8914 | H + SelF-R 0.8871 | H + MR 0.8840 | SelF-R 0.8832 | MultiR 0.8814 | H-IT 0.8761 | MiniR 0.8744 | QUANT 0.8671 |
+| HC2 0.8914 | Mean-Difference rc / r=c / rc Wilcoxon p-value | 0.0043 60 / 3 / 49 0.2374 | 0.0074 59 / 7 / 46 0.1046 | 0.0082 73 / 3 / 36 0.0001 | 0.0100 67 / 6 / 39 0.0051 | 0.0153 73 / 4 / 35 0.0007 | 0.0170 84 / 5 / 23 1e-04 | 0.0243 81 / 5 / 26 1e-04 |
+| H + SelF-R 0.8871 | -0.0043 49 / 3 / 60 0.2374 | - | 0.0031 56 / 4 / 52 0.4733 | 0.0039 75 / 4 / 33 1e-04 | 0.0056 64 / 4 / 44 0.0212 | 0.0110 68 / 3 / 41 0.0061 | 0.0127 87 / 4 / 21 1e-04 | 0.0200 77 / 4 / 31 1e-04 |
+| H + MR 0.8840 | -0.0074 46 / 7 / 59 0.1046 | -0.0031 52 / 4 / 56 0.4733 | - | 0.0008 67 / 5 / 40 0.0015 | 0.0026 68 / 9 / 35 0.0003 | 0.0079 66 / 4 / 42 0.0644 | 0.0096 80 / 5 / 27 1e-04 | 0.0169 75 / 5 / 32 1e-04 |
+| SelF-R 0.8832 | -0.0082 36 / 3 / 73 0.0001 | -0.0039 33 / 4 / 75 1e-04 | -0.0008 40 / 5 / 67 0.0015 | - | 0.0017 47 / 5 / 60 0.1561 | 0.0071 62 / 5 / 45 0.2066 | 0.0088 78 / 7 / 27 1e-04 | 0.0161 74 / 5 / 33 0.0001 |
+| MultiR 0.8814 | -0.0100 39 / 6 / 67 0.0051 | -0.0056 44 / 4 / 64 0.0212 | -0.0026 35 / 9 / 68 0.0003 | -0.0017 60 / 5 / 47 0.1561 | - | 0.0054 63 / 4 / 45 0.1944 | 0.0070 80 / 5 / 27 1e-04 | 0.0143 75 / 5 / 32 1e-04 |
+| H-IT 0.8761 | -0.0153 35 / 4 / 73 0.0007 | -0.0110 41 / 3 / 68 0.0061 | -0.0079 42 / 4 / 66 0.0644 | -0.0071 45 / 5 / 62 0.2066 | -0.0054 45 / 4 / 63 0.1944 | - | 0.0017 58 / 4 / 50 0.5043 | 0.0090 61 / 5 / 46 0.0502 |
+| MiniR 0.8744 | -0.0170 23 / 5 / 84 1e-04 | -0.0127 21 / 4 / 87 1e-04 | -0.0096 27 / 5 / 80 1e-04 | -0.0088 27 / 7 / 78 1e-04 | -0.0070 27 / 5 / 80 1e-04 | -0.0017 50 / 4 / 58 0.5043 | - | 0.0073 60 / 5 / 47 0.0953 |
+| QUANT 0.8671 | -0.0243 26 / 5 / 81 1e-04 | -0.0200 31 / 4 / 77 1e-04 | -0.0169 32 / 5 / 75 1e-04 | -0.0161 33 / 5 / 74 0.0001 | -0.0143 32 / 5 / 75 1e-04 | -0.0090 46 / 5 / 61 0.0502 | -0.0073 47 / 5 / 60 0.0953 | If in bold, then p-value 0.05 |
+The average accuracy rank of these classifiers is displayed in Figure 5.
+In terms of performance, Hydra + SelF-Rocket ranks second only to HIVE-COTE v2.0, and outperforms others.
+SelF-Rocket ranks just after MultiRocket.
+Figure 6 depicts the Multiple Comparison Matrix [22], i.e.
 the full pairwise comparison between the best performing methods tested.
-The mean difference in accuracy between SelF-Rocket and its baseline MiniRocket is approximately 1.23 %
+The mean difference in accuracy between SelF-Rocket and its baseline MiniRocket is approximately 0.88 %.
+We can also see that SelF-Rocket has a slightly higher accuracy than MultiRocket, despite having fewer wins by pairwise comparison.
 Figure 7 displays the mean accuracy, the pairwise win/draw/loss, and value for statistical tests between the tested methods over 30 resamples for the 112 selected UCR datasets.
-Each point represents a dataset, the more a point is distant to the diagonal line, the more one of the two methods performs better than the other.
-In some cases, SelF-Rocket outperforms Hydra-MultiRocket, but in general, both methods offer comparable performance.
-HIVE-COTE v2.0 also has the advantage over SelF-Rocket in certain datasets.
-There is no significant statistical difference by Wilcoxon signed rank test between SelF-Rocket, Hydra-MultiRocket and HIVE-COTE v2.0
-SelF-Rocket (9,996 or 19,992 features) is faster in the training and prediction phase than MultiRocket and Hydra-MultiRocket due to a lower number of features (resp 50,000 and 50,000 + number of features of Hydra).
+Each point represents a dataset; the more a point is distant from the diagonal line, the more one of the two methods performs better than the other.
+In some cases, Hydra + SelF-Rocket outperforms Hydra + MultiRocket, but in general, both methods offer comparable performance.
+HIVE-COTE v2.0 also has the advantage over Hydra + SelF-Rocket in certain datasets.
+There is no significant statistical difference by the Wilcoxon signed rank test between Hydra + SelF-Rocket, Hydra + MultiRocket, and HIVE-COTE v2.0.
+SelF-Rocket (9,996 or 19,992 features) and Hydra + SelF-Rocket (9,996 or 19,992 features + number of features of Hydra) are faster in the classifier prediction phase than MultiRocket (50,000 features) and Hydra + MultiRocket (50,000 + number of features of Hydra) due to a lower number of features.
+Figure 8 shows the IR-PO combinations selected from the 30 resamples in each of the 112 UCR data sets in three scenarios: before the vote validation step (cf Figure 4 and Algorithm 2), when only votes with high consensus are considered (at least 90% of voters must include the chosen combination in their top 5), and after the vote validation step (SelF-Rocket method).
+For the latter, we can see that several IR-PO combinations can be chosen, even if the PPV_MIX combination remains the most frequently chosen, mainly due to its use as the default combination when insufficient consensus is reached (cf Algorithm 2).
+Before the vote validation phase, several different IR-PO combinations could also be selected.
+However, experiments (cf 5.3.5) show suboptimal performance, hence the need for a vote validation phase to ensure the selection of a better IR-PO combination.
+Finally, the distribution of IR-PO combinations for votes resulting in high voter consensus (votes requiring readjustment are not considered here) is closest to the distribution of the best Oracle IR-PO combinations (cf Figure 1).
 5.3 Sensitivity Analysis and Ablative Study
-As previously shown, SelF-Rocket outperforms other existing methods based on random convolution kernels.
-In this section, the impact of the three key parameters on global accuracy performance is studied:
-Following [21, 19], we choose 50 ‘development’ datasets from the 112 UCR datasets to conduct this analysis.
-We first remove the 12 datasets with the most training examples, then we select the sets that contain at least 5 examples per class and 100 training examples.
-From the remaining list, 50 were randomly selected.
+In this section, we examine the influence of the five key parameters of the SelF-Rocket method on overall accuracy performance:
+Following [16, 15], we choose 55 ‘development’ datasets from the 112 UCR datasets to conduct this analysis.
+We first select the set of datasets that contain at least 5 examples per class and 100 training examples.
+From that remaining list, 55 datasets were randomly selected.
 The complete list is available in Table 4.
-The others parameters of SelF-Rocket, i.e. number of kernels, padding, dilation, bias, length of kernel, values inside the kernel, remain the same as the default parameters of MiniRocket.
+5.3.1 Input Representations and Pooling Operators
+| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
+| ONLY BASE | ONLY DIFF | ONLY MIX | ||
+| Win | – | 17 | 16 | 32 |
+| Draw | – | 7 | 3 | 6 |
+| Lose | – | 31 | 36 | 17 |
+| Mean accuracy | 85.15 | 84.61 | 82.51 | 85.16 |
+We evaluated the classification performance of the features generated by 3 sets of input representations: the raw time series (}), the first-order difference applied to the time series (}), and a concatenation of features for both (}).
+Figure 9 illustrates the discrepancy in ranking between the utilization of all three sets of representations and the exclusive use of a single one.
+Table 2 provides a summary of the pairwise win/draw/loss between default SelF-Rocket (FULL) and SelF-Rocket with only one set of representations.
+The first-order difference (DIFF) is generally the least impactful representation, while the MIX set of representations demonstrates the best overall performance among the three and is similar to the FULL version in terms of performance.
+| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
+| NO PPV | NO GMP | NO MIPV | NO MPV | NO LSPV | ||
+| Win | – | 19 | 7 | 15 | 17 | 21 |
+| Draw | – | 6 | 46 | 29 | 16 | 24 |
+| Lose | – | 30 | 2 | 11 | 22 | 10 |
+| Mean acc. | 85.15 | 85.01 | 85.12 | 84.94 | 85.07 | 85.16 |
+The choice of a pooling operator or a set of pooling operators to generate features can affect the performance of the classifier in the classification process.
+Figure 10 and Table 3 emphasize how not employing one of the five pooling operators affects performance.
+It can be seen that not using GMP or LSPV has no significant impact on performance, whereas not using PPV has a greater impact on overall performance.
 5.3.2 Number of folds
-The SelF-Rocket Feature Selection module employs a stratified -fold decomposition of the training set.
-We have conducted tests using the minimal number of folds possible, namely , up to a fixed value of .
-Figure 7 illustrates the impact of the number of folds on the classifier mean accuracy.
-No notable differences were observed when different values of were used.
+[…]
 5.3.3 Number of features
@@
 The number of features tested ranges from 2,500 to 9,996 for each mini-classifier within the Feature Selection Module, features are randomly selected from the total number of features, which is either 9,996 or 19,992 ().
-Figure 7 shows that the number of features has a relatively minor impact on the average accuracy of the final classifier.
-5.3.3 Input Representations and Pooling Operators
-| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
-| ONLY BASE | ONLY DIFF | ONLY MIX | ||
-| Win | – | 2 | 4 | 6 |
-| Draw | – | 18 | 15 | 23 |
-| Lose | – | 30 | 31 | 21 |
-| Mean accuracy | 85.69 | 85.01 | 82.13 | 85.35 |
-We evaluated the classification performance of the features generated by 3 sets of input representations: the raw time series (}), the first-order difference applied to the time series (}), and a concatenation of features for both (}).
-Figure 8 illustrates the discrepancy in ranking between the utilization of all three sets of representations and the exclusive use of a single one.
-Table 2 provides a summary of the pairwise win/draw/loss between SelF-Rocket () and SelF-Rocket with only one set of representations.
-The first-order difference is generally the least impactful representation, while the MIX set of representations demonstrates the best overall performance among the three.
-| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
-| NO PPV | NO GMP | NO MIPV | NO MPV | NO LSPV | ||
-| Win | – | 4 | 1 | 1 | 0 | 4 |
-| Draw | – | 27 | 48 | 39 | 46 | 43 |
-| Lose | – | 19 | 1 | 10 | 4 | 3 |
-| Mean acc. | 85.69 | 85.44 | 85.63 | 85.36 | 85.60 | 85.67 |
-The choice of a pooling operator or a set of pooling operators to generate features can affect the performance of the classifier in the classification process.
-Figure 9 and Table 3 show the impact on performance of not using one of the five pooling operators.
-It can be seen that not using GMP or LSPV has no significant impact on system performance.
-On the other hand, not using MIPV or PPV has a greater impact on overall performance.
+As exposed in Figure 12, the number of features has a relatively minor impact on the average accuracy of the final classifier.
+5.3.4 Number of runs
+The total number of decompositions of the initial data set depends on both the number of folds and the number of runs .
+The first one varies the distribution of the number of examples between the training set and the validation set, while the second one allows the original training set to be shuffled and then re-split differently.
+When , Figure 12 shows that the choice of the number of runs is relatively more important to improve the performance of the Features Selection Module.
+5.3.5 Vote Validation
+Once the highest median vote has been carried out, using the result of the vote directly can be risky, especially for small datasets because their mini-classifiers are trained on datasets that are too small when selecting the best set of features.
+To avoid selecting a wrong combination that would generalize poorly, a vote validation system has been put in place by calculating the percentage of voters, who more or less agree with our final choice.
+If this percentage exceeds a certain value (threshold in Algorithm 2) then the vote is maintained, otherwise the vote is replaced by a default choice (PPV_MIX).
+Figure 13 displays the variation of the mean accuracy given a threshold value and a , e.g.
+if the , we calculate the percentage of voters with the final choice in their 4 best accuracy.
+In Figure 13, we can see that the higher the top, the higher the threshold thresh required to obtain greater accuracy.
+However, if the threshold is too high, the IR-PO combination selected is replaced too often by the default combination, which gives a lower accuracy.
+A slight improvement in performance can be observed with the use of the vote validation system, but a greater impact of this system can be observed with smaller datasets cf. Appendix B.
 5.4 Scalability and Compute Time Comparison
-Compared to other ROCKET methods, SelF-Rocket incorporates a Feature Selection module between the global set of features and the final classifier.
-The time required for the generation of SelF-Rocket features is less than Hydra-MultiRocket, while the training of the classifier is similar in duration to that of MiniRocket.
-Figure 10 shows the mean compute time for all resamples and all datasets of the Feature Selection module according to the number of folds and features.
+Compared to other ROCKET methods, SelF-Rocket includes a Feature Selection Module between the global feature set and the final classifier.
+The feature generation time of SelF-Rocket and Hydra + SelF-Rocket is similar to MultiRocket and Hydra + MultiRocket respectively, while the training time of SelF-Rocket’s Ridge classifier is similar to that of MiniRocket.
+Figure 14 shows the mean computation time for all resamples and all datasets of the Feature Selection Module according to the number of runs and features.
 The classifier complexity depends on , the number of features and , the number of training examples.
-is a multiplier for the number of mini-classifiers, and impacts the complexity of a single classifier so the two parameters impact generally the compute time.
+The number of runs is a multiplier for the number of mini-classifiers, and impacts the complexity of a single classifier so the two parameters impact generally the compute time.
+The comparison of the performance of SelF-Rocket (FULL) and SelF-Rocket (MIX) of Section 5.3.1 reveals that they are comparable.
+However, the feature selection phase of SelF-Rocket (MIX) is faster due to a lower number of IR-PO combinations (only 5).
+With , and , the SelF-Rocket Feature Selection Module takes 3.71 seconds on average for all resamples with the MIX version, compared to 11.12 seconds on average for the FULL version.
 For each combination of IR and PO, we instantiate mini-classifiers to identify the optimal set of features.
 A Ridge classifier with a complexity of is employed.
-In Figure 11(a) given that is a fixed value, namely 9,996, we can see that the mean time only depends on the number of training examples.
-Figure 12(a) shows the average total computation time over the 50 development datasets, for all resamples.
-SelF-Rocket and MiniRocket training (Figure 12(a)) and prediction times (Figure 12(b)) are comparable.
+In Figure 15(a) given that is a fixed value, namely 5,000, we can see that the mean time only depends on the number of training examples.
+We have a value of (cf Algorithm 1), which explains why there is a cap on the Feature Selection Module’s time after this value.
+In Figure 16(a), we can see the average total computation time over the 55 development datasets, for all resamples for the ROCKET methods.
+SelF-Rocket and MiniRocket Ridge classifier’s training (Figure 16(a)) and prediction times (Figure 16(b)) are comparable.
 However, among the evaluated methods, MiniRocket remains the fastest overall.
-As illustrated in Figure 10, with and SelF-Rocket’s average feature selection time is around 1.3 seconds, which is significantly faster than its feature creation computation time.
 SelF-Rocket features take longer to create than MultiRocket features.
@@
 This explains why HYDRA is the fastest in terms of classifier training and testing time.
-As illustrated in Figures 12(c) and 12(d), SelF-Rocket is faster than Hydra-MultiRocket in terms of both feature creation time and classifier testing time, with less discrepancy.
+As illustrated in Figures 16(c) and 16(d), Hydra + SelF-Rocket and SelF-Rocket are both faster than Hydra + MultiRocket and MultiRocket in terms of classifier testing time, with less discrepancy.
+5.5 SelF-Rocket with an oracle
+As shown in Section 5, selecting the best features can be difficult, particularly in small datasets.
+Therefore, the implementation of SelF-Rocket presented in this paper is not optimal.
+Knowing the optimal input representations and pooling operator, what would be the performances of SelF-Rocket?
+Figure 17(a) illustrates the average mean accuracy ranking for the various classifiers previously tested, while Figure 17(b) presents the MCM of the Oracle version of SelF-Rocket and Hydra combined with SelF-Rocket, compared to HIVE-COTE v2.0 and other ROCKET methods.
+These results indicate the maximum potential performance of an algorithm based on SelF-Rocket.
+The question remains whether such performance is attainable, or if we can approach it with an improved Feature Selection Module.
+| Mean-Accuracy | H + SelF-R (Oracle) 0.8916 | HC2 0.8914 | SelF-R (Oracle) 0.8892 | H + MR 0.8840 | MultiR 0.8814 | MiniR 0.8744 |
+| H + SelF-R (Oracle) 0.8916 | Mean-Difference rc / r=c / rc Wilcoxon p-value | 0.0002 66 / 5 / 41 0.0236 | 0.0024 68 / 6 / 38 0.0008 | 0.0076 72 / 8 / 32 1e-04 | 0.0101 83 / 8 / 21 1e-04 | 0.0172 97 / 5 / 10 1e-04 |
+| HC2 0.8914 | -0.0002 41 / 5 / 66 0.0236 | - | 0.0023 57 / 5 / 50 0.6579 | 0.0074 59 / 7 / 46 0.1046 | 0.0100 67 / 6 / 39 0.0051 | 0.0170 84 / 5 / 23 1e-04 |
+| SelF-R (Oracle) 0.8892 | -0.0024 38 / 6 / 68 0.0008 | -0.0023 50 / 5 / 57 0.6579 | - | 0.0052 61 / 8 / 43 0.0261 | 0.0077 69 / 7 / 36 0.0001 | 0.0147 97 / 6 / 9 1e-04 |
+| H + MR 0.8840 | -0.0076 32 / 8 / 72 1e-04 | -0.0074 46 / 7 / 59 0.1046 | -0.0052 43 / 8 / 61 0.0261 | - | 0.0026 68 / 9 / 35 0.0003 | 0.0096 80 / 5 / 27 1e-04 |
+| MultiR 0.8814 | -0.0101 21 / 8 / 83 1e-04 | -0.0100 39 / 6 / 67 0.0051 | -0.0077 36 / 7 / 69 0.0001 | -0.0026 35 / 9 / 68 0.0003 | - | 0.0070 80 / 5 / 27 1e-04 |
+| MiniR 0.8744 | -0.0172 10 / 5 / 97 1e-04 | -0.0170 23 / 5 / 84 1e-04 | -0.0147 9 / 6 / 97 1e-04 | -0.0096 27 / 5 / 80 1e-04 | -0.0070 27 / 5 / 80 1e-04 | If in bold, then p-value 0.05 |
 6 Conclusion and prospects
-The two main contributions of this article are as follows:
-First, we have shown that the choice of Input Representations and Pooling Operators significantly impacts the performance of TSC methods based on random convolution kernels.
-We have also shown that using multiple Pooling Operators, such as MultiRocket, leads to good performances on average, but is not optimal in specific cases: it is always preferable to select a single appropriate Pooling Operator.
-Then, on this basis, we have proposed a new algorithm based on MiniRocket, SelF-Rocket, incorporating a feature selection stage.
-Experiments show that this algorithm selects the best Input Representations / Pooling Operator combination in most cases, leading to state of the art performances.
-In the tested implementation of SelF-Rocket, only 9,996 or 19,992 features are retained to train the RIDGE classifier, far fewer than competing methods such as MultiRocket or Hydra-MultiRocket, leading to fast predictions.
-In this article, a minimal version of SelF-Rocket was evaluated, using 2 input representations, 5 pooling operators and a simple wrapper-based feature selection method.
-This implementation could be improved by:
-Considering more input representations (second order difference, Fourier and Hilbert transforms, etc.),
-Combining SelF-Rocket and HYDRA features,
-Using HYDRA as a pooling operator within SelF-Rocket,
+This article presents a novel algorithm called SelF-Rocket, which builds upon MiniRocket by incorporating a feature selection stage.
+The key idea behind SelF-Rocket is its dynamic selection of the best combination of input representations and pooling operators during the training phase.
+Experiments conducted on the UCR datasets show that SelF-Rocket, on average, selects a more effective combination of Input Representation and Pooling Operator compared to MiniRocket, and succeeds to achieve state-of-the-art performance.
+SelF-Rocket delivers performance comparable to MultiRocket (similarly, Hydra combined with SelF-Rocket matches the performance of Hydra combined with MultiRocket).
+However, SelF-Rocket offers the benefit of quicker classification prediction times due to its utilization of fewer features (9,996 or 19,992 for SelF-Rocket vs 50,000 for MultiRocket).
+The version of SelF-Rocket evaluated in this study employs three sets of input representations, five pooling operators, and a simple wrapper-based feature selection method.
+Potential improvements to this implementation could include:
+Incorporating additional input representations, such as second-order differences, Fourier transforms, and Hilbert transforms, among others.
+Improving the vote validation systems could potentially bring performance closer to the optimal levels achieved by the oracle.
 Using a filter-based feature selection method (based on test for instance) instead of Stratified -Fold to speed-up the feature selection stage.
@@
 This research project, supported and financed by the French ANR (Agence Nationale pour la Recherche), is part of the Labcom (Laboratoire Commun) MYEL (MobilitY and Reliability of Electrical chain Lab) involving LSEE, LGI2A and CRITTM2A (ANR-22-LCV2-0001 MYEL).
+The authors would like to thank Prof.
+Eamonn Keogh and all those who have contributed, and continue to contribute, to the maintenance of the University of California Riverside (UCR) TSC benchmark datasets and to open source implementations of the algorithms used in this article:
+The original implementations of MiniRocket555https://github.com/angus924/minirocket, MultiRocket666https://github.com/ChangWeiTan/MultiRocket and Hydra777https://github.com/angus924/hydra were used as baseline.
+Critical Difference diagrams were generated using aeon toolkit [23].
+Multiple Comparison Matrices were generated using the original implementation888https://github.com/MSD-IRIMAS/Multi_Comparison_Matrix of [18].
+The results provided by tsml-eval999https://github.com/time-series-machine-learning/tsml-eval were used to benchmark SelF-Rocket.
 Appendix A Additional Figures & Tables
-| Developement datasets | ||
+| Development datasets | ||
 | ACSF1 | ChlorineConcentration | Computers |
-| CricketX | CricketY | CricketZ |
-| DistalPhalanxOutlineAgeGroup | DistalPhalanxOutlineCorrect | DistalPhalanxTW |
-| Earthquakes | ECG200 | EOGHorizontalSignal |
-| EOGVerticalSignal | EthanolLevel | FaceAll |
-| Fish | FreezerRegularTrain | GunPointAgeSpan |
+| CricketX | CricketY | DistalPhalanxOutlineAgeGroup |
+| DistalPhalanxOutlineCorrect | DistalPhalanxTW | Earthquakes |
+| ECG200 | EOGHorizontalSignal | EOGVerticalSignal |
+| EthanolLevel | FaceAll | Fish |
+| FordA | FreezerRegularTrain | GunPointAgeSpan |
 | GunPointMaleVersusFemale | GunPointOldVersusYoung | Ham |
@@
 | LargeKitchenAppliances | MedicalImages | MiddlePhalanxOutlineAgeGroup |
-| MiddlePhalanxOutlineCorrect | MixedShapesRegularTrain | MixedShapesSmallTrain |
-| OSULeaf | Plane | PowerCons |
-| ProximalPhalanxOutlineAgeGroup | ProximalPhalanxOutlineCorrect | ProximalPhalanxTW |
-| ScreenType | SemgHandGenderCh2 | SemgHandMovementCh2 |
-| SemgHandSubjectCh2 | ShapesAll | SmoothSubspace |
-| SwedishLeaf | SyntheticControl | Trace |
-| UWaveGestureLibraryX | UWaveGestureLibraryY | UWaveGestureLibraryZ |
-| Worms | WormsTwoClass | |
+| MiddlePhalanxOutlineCorrect | MiddlePhalanxTW | MixedShapesRegularTrain |
+| MixedShapesSmallTrain | OSULeaf | PhalangesOutlinesCorrect |
+| Plane | PowerCons | ProximalPhalanxOutlineAgeGroup |
+| ProximalPhalanxOutlineCorrect | ProximalPhalanxTW | RefrigerationDevices |
+| ScreenType | SemgHandMovementCh2 | SemgHandSubjectCh2 |
+| ShapesAll | SmallKitchenAppliances | SmoothSubspace |
+| StarLightCurves | Strawberry | SwedishLeaf |
+| SyntheticControl | Trace | UWaveGestureLibraryX |
+| UWaveGestureLibraryZ | Worms | WormsTwoClass |
+| Yoga | ||
+Appendix B Small vs Large datasets
+We use all of the 36 UCR datasets with less than 100 training examples (thus missing from the sensitivity analysis) to see the impact of the number of runs 5.3.4 and vote validation 5.3.5 on small datasets.
+Figure 20 shows that these 2 SelF-Rocket components have an even greater effect on small datasets than on larger datasets.
+Regardless of the top considered, a significant consensus among voters is necessary to achieve optimal performance for the smallest datasets, as outlined in Figure 20(a).
+As Figure 20(b) illustrates it, using these two components together gets better results than using them separately.
+The time needed to train SelF-Rocket depends on the training size.
+The larger the training size, the less important the time for SelF-Rocket feature selection becomes compared to the other time, as shown, for example, in Figure 21.